OpenStudy (anonymous):

the sum of 6 and -12 decreased b y 21 write an algebraic expression and simplify

OpenStudy (bahrom7893):

(6-12)-21=-27

OpenStudy (anonymous):

I can always count on U thNX

OpenStudy (bahrom7893):

np

OpenStudy (anonymous):

7/4\[\div\] - 21/2

OpenStudy (anonymous):

why would (-21/2) have brackets like this?

OpenStudy (bahrom7893):

because it's negative.. but it doesn't really matter..

OpenStudy (bahrom7893):

they usually put parenthesis around negative numbers..

OpenStudy (anonymous):

but in one case there positive absolute value right? thats just lines then?

OpenStudy (anonymous):

ok.

OpenStudy (bahrom7893):

yes if u have absolute value then u just have lines.. | |

OpenStudy (anonymous):

ok thank you and if then theres a negative number on outside it remains negative

OpenStudy (anonymous):

so that prob threw me off I got it thanx

OpenStudy (bahrom7893):

okay np.. -1/6 is the answer just in case..

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

how did u determine it was negative?

OpenStudy (bahrom7893):

well the negative from -21/2 doesn't go anywhere.. unless u multiply it by another negative..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(\[\div\]

OpenStudy (anonymous):

8\[(8\div 4)^{2}\]

OpenStudy (anonymous):

forget about the eighht above that im not sure how to do that though

OpenStudy (bahrom7893):

8 divided by 4 is 2... 2 squared is equal to 2*2=4 (8/4)^2 = (2)^2=4

OpenStudy (anonymous):

you square the 2 ok

OpenStudy (anonymous):

-8/11+ 1/2

OpenStudy (bahrom7893):

sorry post this as new thread.. i won't be able to help u i gtg and nobody else will notice if u keep posting it here..